# Latent Debate: A Surrogate Framework for Interpreting LLM Thinking

### v1

@@ Abstract
Understanding the internal thinking process of Large Language Models (LLMs) and the cause of hallucinations remains a key challenge. To this end, we introduce latent debate, a novel framework for interpreting model predictions through the lens of implicit internal arguments. Unlike the current work of self-consistency and multi-agent debate, which relies on explicit debates among multiple answers or multiple models, latent debate captures the hidden supporting and attacking signals that arise within a single model during a single inference. We first present a model- and task-agnostic conceptual framework, and then instantiate it symbolically to approximate the thinking process of LLMs on True/False prediction tasks. Empirical studies demonstrate that latent debate is a faithful structured surrogate model that has highly consistent predictions with the original LLM. Beyond interpretability, we demonstrate that latent debate provides a strong baseline for hallucination detection. Further analysis reveals strong correlations between hallucinations and debate patterns, such as a high degree of latent debates in the middle layers is linked to a higher risk of hallucinations. These findings position latent debate as a potential framework for understanding internal mechanisms of LLMs, especially for scenarios where internal (dis)agreements appear during the inference steps. The code is available at   https://github.com/tigerchen52/latent_debate
@@ 1 Introduction
Large Language Models (LLMs) have made remarkable progress on many reasoning tasks, yet they continue to suffer from hallucinations (xu2024hallucination; huang2025survey). For example, LLMs may generate answers that contradict user prompts and or conflict with the source of training data (ji2023survey; kalai2025why; bang2025hallulens), seriously undermining their reliability and trustworthiness. This is further aggravated by the fact that, due to their opacity, it is difficult to understand why LLMs make given predictions, or why their “thinking” process is flawed.
Recent work in mechanistic interpretability has examined hallucinations through various internal signals, including activations (ferrando2025know), attention patterns (chuang2024lookback), and hidden states (azaria-mitchell-2023-internal). Another line of relevant research (wang2022self) leverages external consistency, i.e., the agreement among multiple answers, to analyze hallucination behaviors. Their findings reveal that hallucinated outputs tend to have low self-consistency (wang2022self). This phenomenon suggests that strong agreement among multiple answers often yields more certain and accurate answers, whereas disagreement indicates higher uncertainty and can serve as a good signal for understanding hallucinations. Subsequent approaches (irving2018ai; du2024improving; chen2024reconcile; liang2024encouraging) further introduce Multiple-Agent Debate (MAD) to reduce hallucinated answers via a debate process of multiple language models, often outperforming single-model baselines.
Inspired by these prior studies of mechanistic interpretability and disagreement/debate, we aim to understand how hallucinations emerge within a model but shifting to latent debate, i.e., arising among different layers and “thinking” steps within an individual model and a single inference  (chuang2023dola; liang2024internal; xie2024calibrating), rather than externally to it as in prior work. Unlike in a conventional, human debate, where arguments are in natural language, the arguments in our latent debate are to be understood in a metaphorical sense as they correspond to latent states in models. Intuitively, latent debates aggregate metaphorical arguments to reflect the thinking and decision-making process beneath the surface. In psychological theories, the human thinking process often involves internal debate-like behaviors such as inner speech (barker1966model) and the dialogical self (hermans2001dialogical). Here, we extend this psychological insight to models introducing latent debates to describe an analogous process taking place within a model.
We focus on two key research questions in this work: (1) Can we use latent debate to model the LLM thinking process? (2) Can latent debate identify hallucinations?
((a))
((b))
@@
To answer the first question, we present a conceptual framework of latent debate that depicts (dis)agreement within a model, which is model- and task-agnostic. The framework consists of three abstract components: latent arguments derived from internal signals, an argument interpreter that translates these implicit arguments into human-readable opinions such as supporting or attacking a claim, and a thinking module that aggregates them to make the final decisions. We then instantiate this framework in decoder-based LLMs on True/False prediction tasks (see case studies in Figure 1), where hidden states serve as latent arguments, the unembedding matrix acts as the argument interpreter, and the thinking module is realized through a symbolic argumentation framework, in the spirit of (vcyras2021argumentative). An empirical study demonstrates that this latent debate acts as a structured surrogate model, providing a faithful approximation of LLM thinking, which achieves up to 98.3% consistency with LLaMA-13B decisions. These findings validate that our latent debate can imitate the thinking process of LLM true/false tasks.
@@
To address the second question, we extract features from the latent debate graph, e.g., the number of internal debates and argument strengths, and train a simple MLP classifier to distinguish hallucinated from non-hallucinated outputs. We find that our latent debate can achieve highly competitive performance in hallucination detection. We then use SHAP attribution scores (SHAP) to identify which features most strongly drive hallucination predictions. Our analyses indicate that a high degree of latent debate, particularly in the middle layers, is the strongest predictor of hallucination.
@@
In summary, our contributions are threefold. (1) We propose latent debate, a novel, model-agnostic framework that leverages internal arguments to interpret a model’s thinking process. (2) We present a symbolic instantiation of latent debate that serves as a faithful surrogate for LLM True/False tasks. (3) We develop a debate-based MLP to detect hallucinations, which help us identify distinct debate patterns, especially intense internal debates in middle layers, that correlate strongly with hallucination in LLMs.
@@ 2.1 Multiple Agent Debate
Multiple-agent debate (MAD) has emerged as a powerful approach for improving factuality and reasoning. Pioneering work on AI safety via debate (irving2018ai) models debate as a self-play game with a (human) judge and provides core theoretical motivation. Recent work adopts multiple language model agents to debate over individual responses jointly, with the final decision made either through consensus (du2024improving; chen2024reconcile) or by a judge (liang2024encouraging). This debate strategy can outperform single model baselines on a wide variety of reasoning tasks. Subsequent research proposes refined debate approaches (li2024improving; liu2024groupdebate) using MAD as an evaluator (chanchateval). In this work, we focus on debates operating implicitly within a model or agent rather than externally visible debates among multiple agents.
@@ 2.2 Internal Consistency
Another line of work focuses on how to use the consistency of internal model states, such as logits and activations, to improve model outputs (liang2024internal). For example, DoLa (chuang2023dola) proposes a decoding strategy that contrasts logits between later layers and earlier layers for improving LLM truthfulness. xie2024calibrating adopts internal consistency, i.e., how middle layers’ predictions (dis)agree with the final layer, to guide LLM decoding. In this work, we aim to use latent (dis)agreements to build a surrogate framework for interpreting the thinking process of a model rather than enhancing model outputs.
@@ 2.3 Computational Argumentation in Explainable AI
Argumentation Frameworks (AFs) (Dung_95), are a fundamental formalism in computational argumentation (CA-survey; vcyras2021argumentative), a well-established research area in AI. According to (Dung_95), an AF consists of a set of arguments and a binary attack relation among them. Arguments are seen as abstract entities, while the attack relation captures conflicts between arguments. AFs and their extensions, such as incorporating weights for arguments and a support relation between arguments, collectively referred to as Quantitative Bipolar Argumentation Frameworks (QBAFs) (rago2016discontinuity; baroni2019fine; vcyras2021argumentative), have been widely adopted in Explainable AI (XAI). AFs can serve as surrogate models to approximate the inner structure and decision-making process of AI systems (vcyras2021argumentative; potyka2021interpreting; potyka2023explaining; ayoobi2023sparx). Beyond serving as surrogates, AFs can also be explicitly integrated into AI or LLM systems to enhance explainability(freedman2025argumentative; vcyras2021argumentative; vassiliades2021argumentation; engelmann2022argumentation; guo2023argumentative). In this work, we are the first to introduce the model-agnostic concept of latent debate and adopt an AF as the thinking module of a model.
@@ 3.1 Problem Statement
In this work, we aim to obtain a structured surrogate model to understand the internal ‘thinking’ of a target model, as opposed to a conventional surrogate model solely imitating the input-output behaviour of the model. More specifically, given a target model $M:\mathcal{X}\to\mathcal{Y}$, a conventional surrogate model $S:\mathcal{X}\to\mathcal{Y}$ imitates the input-output behaviour of the target model, i.e., $S(x)\approx M(x)$. Such surrogates offer a simplified and interpretable approximation of the model’s outputs (asher2015review; kudela2022recent).
This type of surrogate does not faithfully reflect the thinking (or decision-making) process of the model. To address this limitation, structured surrogate models explicitly accounts for the model’s internal organization (munk2022probabilistic; paez2024understanding). Supposing the target model has a known structure:
@@
where $h(x)\in\mathcal{H}$ is the internal structure, our goal is to construct:
@@
that approximates the internal computational structure of the target model, i.e., $\hat{h}(x)\approx h(x)$, while, at the same time, being faithful to the target model by making highly consistent predictions with it, i.e., $\hat{y}(x)\approx y(x)$.
@@ 3.2 Conceptual Framework
A latent argument refers to an internal signal within a model that implicitly conveys supporting or attacking opinions toward a claim. Such signals can derive from different sources, like activations or attention patterns. Because they live in the model’s latent space, these arguments are not directly visible or human-readable, but they still express how intermediate steps “think” about the claim.
@@
3.3.1 Instantiation
We now describe how the abstract concepts of latent debate can be instantiated in the context of transformer-based LLM true/false prediction tasks (vaswani2017attention). We adopt a symbolic argumentation framework to perform the decision making process, which is transparent and efficient.
Formally, given a query $\mathbf{x}=(x_{1},\ldots,x_{N})$ with a binary label $c\in\{\text{True},\text{False}\}$, a decoder-based LLM generates an answer $\mathbf{y}=(y_{1},\ldots,y_{T})$. Both the query and answer tokens are drawn from the same vocabulary, i.e., $x_{n},y_{t}\in\mathcal{V}$. Each token $y_{t}$ in the answer is generated conditionally based on the preceding tokens and the input query, following the distribution: $y_{t}\sim P(y_{t}\mid\mathbf{y}_{\leq t-1},\mathbf{x})$. In a MAD setting (du2024improving; liang2024encouraging), the process involves generating multiple answers $\mathcal{Y}=\{\mathbf{y}^{(1)},\ldots,\mathbf{y}^{(K)}\}$, which may support ($\mathbf{y}^{+}$) or attack ($\mathbf{y}^{-}$) the claim. A final decision is then derived by aggregating these arguments, often through some form of consensus or voting strategies. In contrast, we define that a latent debate takes place inside the model processing claim $\mathbf{x}$ before generating answer $\mathbf{y}$.
Latent Arguments in LLMs.
@@ 3.3.1 Instantiation
Argument Interpreter in LLMs.
To make latent arguments interpretable, the instantiated argument interpreter projects a hidden state into the vocabulary space using the unembedding matrix $\mathbf{W}^{\text{unemb}}\in\mathbb{R}^{|\mathcal{V}|\times d}$. This produces a probability distribution over vocabulary tokens, which has been widely used in mechanistic studies (nostalgebraist2020interpretinggpt; belrose2023elicitinglatentpredictionstransformers). By examining the probabilities assigned to specific tokens True and False, we can quantify the opinion of each latent argument, i.e., how much it supports or attacks the claim.
@@
Thinking Module in LLMs.
@@
$\text{cls}(\mathbf{x})=\text{think}(\mathcal{H}),\ \mathcal{H}=\{\text{interpret}(\mathbf{h}_{n}^{(l)})\mid 1\leq n\leq N,\ 1\leq l\leq L-1\}$ (5)
where $\text{cls}(\mathbf{x})$ is capable of outputting a label associated with the final decision $c$. To perform the thinking step, we adopt a symbolic approach, Quantitative Bipolar Argumentation Framework (QBAF) (baroni2019fine), as the $\textit{think}(\cdot)$ function, which accounts for both supporting and attacking relationships among arguments to yield a coherent, weighted judgment.
((a))
((b))
@@
Given a set of arguments $\mathcal{A}$, QBAF is capable of considering the overall debate situation and outputting a final strength (as shown in Figure 3(a)), which can be used to obtain the binary (true/false) predictions. In the figure, each node $n_{i}$ corresponds to an argument, and the $\tau(\cdot)$ function assigns the initial strength, indicating its polarity and strength magnitude before propagation. The directed edges between nodes represent the relationships among arguments: edges labeled with - are attacks, indicating that one argument undermines another (two arguments with different polarities), while edges labeled with ++ are supports, meaning that one argument reinforces another. The $\sigma(\cdot)$ beneath the nodes is a function resulting final strengths after applying a chosen gradual semantics e.g. (baroni2015automatic; rago2016discontinuity; Potyka18; amgoud2018evaluation), which reflects how the collective influence of attackers and supporters modifies the outcome. See Example 1 in the appendix for a detailed computation process.
Formally, gradual semantics provide the rules that control how initial strengths are updated and result in the computation of final strengths. For this goal, we start from arguments with no attackers or supporters, whose final strength is the same as their initial strength. For the remaining arguments, the final strength is updated along the edges by considering the influence of both attackers and supporters. This process involves two components: aggregation and influence. For an argument $\alpha\!\!\in\!\!\mathcal{A}$, the aggregation step computes its energy $E_{\alpha}\!$ by summing the strengths of its attackers and supporters ($\beta$):

### v2

@@ Abstract
Understanding the internal thinking process of Large Language Models (LLMs) and the cause of hallucinations remains a key challenge. To this end, we introduce latent debate, a structured surrogate framework for interpreting model predictions through the lens of implicit internal arguments. Unlike traditional explicit debates, latent debate captures the hidden supporting and attacking signals that arise within a single model during a single inference. We first present a model- and task-agnostic conceptual framework, and then instantiate it symbolically to approximate the thinking process of LLMs on True/False prediction tasks. Empirical studies demonstrate that our latent debate is a faithful structured surrogate model that has highly consistent predictions with the original LLM. Beyond interpretability, we demonstrate that the latent debate provides a strong baseline for hallucination detection. Further analysis reveals strong correlations between hallucinations and debate patterns, such as a high degree of latent debates in the middle layers is linked to a higher risk of hallucinations. These findings position latent debates as a potential framework for understanding internal mechanisms of LLMs, especially for scenarios where internal (dis)agreements appear during the inference steps.
Machine Learning, ICML
@@ 1 Introduction
Large Language Models (LLMs) have made remarkable progress on many reasoning tasks, yet they continue to suffer from hallucinations (Xu et al., 2024; Huang et al., 2025). For example, LLMs may generate answers that contradict user prompts or conflict with the source of training data (Ji et al., 2023; Kalai et al., 2025; Bang et al., 2025), seriously undermining their reliability and trustworthiness. This is further aggravated by the fact that, due to their opacity, it is difficult to understand why LLMs make given predictions, or why their “thinking” process is flawed.
@@
Recent work in mechanistic interpretability has examined hallucinations through various internal signals, including activations (Ferrando et al., 2025), attention patterns (Chuang et al., 2024), and hidden states (Azaria and Mitchell, 2023). Another line of relevant research (Wang et al., 2022) leverages external consistency, i.e., the agreement among multiple answers, to analyze hallucination behaviors. Their findings reveal that hallucinated outputs tend to have low self-consistency (Wang et al., 2022). This phenomenon suggests that strong agreement among multiple answers often yields more certain and accurate answers, whereas disagreement indicates higher uncertainty and can serve as a good signal for detecting hallucinations. Subsequent approaches (Irving et al., 2018; Du et al., 2024; Chen et al., 2024; Liang et al., 2024a) further introduce Multiple-Agent Debate (MAD) to reduce hallucinated answers via a debate process of multiple language models, often outperforming single-model baselines.
Inspired by these prior studies of mechanistic interpretability and disagreement/debate, we aim to understand how hallucinations emerge within a model, but shifting to latent debate, i.e., arising among different layers and “thinking” steps within an individual model and a single inference  (Chuang et al., 2023; Liang et al., 2024b; Xie et al., 2024), rather than externally to it as in prior work. Unlike in a conventional human debate, where arguments are in natural language, the arguments in our latent debate are to be understood in a metaphorical sense as they correspond to latent states in models. Intuitively, latent debates aggregate metaphorical arguments to reflect the thinking and decision-making process beneath the surface. In psychological theories, the human thinking process often involves internal debate-like behaviors such as inner speech (Barker and Wiseman, 1966) and the dialogical self (Hermans, 2001). Here, we extend this psychological insights to models. introducing latent debates to describe an analogous process taking place within a model.
We focus on two key research questions in this work: (1) Can we use latent debate to model the LLM thinking process? (2) Can latent debate identify and interpret hallucinations? We use the concept of latent debate to build a transparent and simple surrogate model that approximates the predictions of a complex black-box model, with the goal of understanding and interpreting the target model’s behavior. Our approach is not intended to improve the target model’s capabilities, such as improving factuality, but rather to provide insights into its internal decision-making process.
@@
To answer the first question, we present a conceptual framework of latent debate that depicts (dis)agreements within a model, while being model- and task-agnostic. The framework consists of three abstract components: latent arguments derived from internal signals, an argument interpreter that translates these implicit arguments into human-readable opinions such as supporting or attacking a claim, and a thinking module that aggregates them to make the final decisions. We then instantiate this framework in decoder-based LLMs on True/False prediction tasks (see case studies in Figure 1), where hidden states serve as latent arguments, the unembedding matrix acts as the argument interpreter, and the thinking module is realized through a symbolic argumentation framework, in the spirit of (Čyras et al., 2021). An empirical study demonstrates that this latent debate acts as a structured surrogate model, providing a faithful approximation of LLM thinking, which achieves up to 97% consistency with LLaMA-13B decisions. These findings validate that our latent debate can imitate the thinking process of LLM true/false tasks.
@@
To address the second question, we extract features from the latent debate graph, e.g., the number of internal debates and argument strengths, and train a simple tree classifier to distinguish hallucinated from non-hallucinated outputs. We find that our latent debate can achieve highly competitive performance in hallucination detection. We then use SHAP attribution scores (Lundberg and Lee, 2017b) to identify which features most strongly drive hallucination predictions. Our analyses indicate that a high degree of latent debate, particularly in the middle layers, is the strongest predictor of hallucination.
@@
In summary, our contribution is threefold. (1) We propose latent debate, a novel, model-agnostic framework that leverages internal arguments to interpret a model’s thinking process. (2) We present a symbolic instantiation of latent debate that serves as a faithful surrogate for LLM True/False tasks. (3) We develop a simple debate-based approach to detect hallucinations, which help us identify distinct debate patterns.
@@ 2.1 Multiple Agent Debate
Multiple-agent debate (MAD) has emerged as a powerful approach for improving factuality and reasoning. Pioneering work on AI safety via debate (Irving et al., 2018) models debate as a self-play game with a (human) judge and provides core theoretical motivation. Recent work adopts multiple language model agents to debate over individual responses jointly, with the final decision made either through consensus (Du et al., 2024; Chen et al., 2024) or by a judge (Liang et al., 2024a). This debate strategy can outperform single model baselines on a wide variety of reasoning tasks. Subsequent research proposes refined debate approaches (Li et al., 2024; Liu et al., 2024) using MAD as an evaluator (Chan et al., 2024). In this work, we focus on debates operating implicitly within a model or agent rather than externally visible debates among multiple agents.
@@ 2.2 Internal Consistency
Another line of work focuses on how to use the consistency of internal model states, such as logits and activations, to improve model outputs (Liang et al., 2024b). For example, DoLa (Chuang et al., 2023) proposes a decoding strategy that contrasts logits between later layers and earlier layers for improving LLM truthfulness. Xie et al. (2024) adopts internal consistency, i.e., how middle layers’ predictions (dis)agree with the final layer, to guide LLM decoding. In this work, we aim to use latent (dis)agreements to build a surrogate framework for interpreting the thinking process of a model rather than enhancing model outputs. We also show that our surrogate model can help detect hallucinations.
@@ 2.3 Computational Argumentation in Explainable AI
Argumentation Frameworks (AFs) (Dung, 1995), are a fundamental formalism in computational argumentation (Atkinson et al., 2017; Čyras et al., 2021), a well-established research area in AI. According to (Dung, 1995), an AF consists of a set of arguments and a binary attack relation among them. Arguments are seen as abstract entities, while the attack relation captures conflicts between arguments. AFs and their extensions, such as incorporating weights for arguments and a support relation between arguments, collectively referred to as Quantitative Bipolar Argumentation Frameworks (QBAFs) (Rago et al., 2016; Baroni et al., 2019; Čyras et al., 2021), have been widely adopted in Explainable AI (XAI). AFs can serve as surrogate models to approximate the inner structure and decision-making process of AI systems (Čyras et al., 2021; Potyka, 2021; Potyka et al., 2023; Ayoobi et al., 2023). Beyond serving as surrogates, AFs can also be explicitly integrated into AI or LLM systems to enhance explainability(Freedman et al., 2025; Čyras et al., 2021; Vassiliades et al., 2021; Engelmann et al., 2022; Guo et al., 2023). In this work, we are the first to introduce the model-agnostic concept of latent debate and adopt an AF as the thinking module of a model.
@@ 3.1 Problem Statement
In this work, we aim to obtain a structured surrogate model to understand the internal ‘thinking’ of a target model, as opposed to a conventional surrogate model solely imitating the input-output behaviour of the model. More specifically, given a target model $M:\mathcal{X}\to\mathcal{Y}$, a conventional surrogate model $S:\mathcal{X}\to\mathcal{Y}$ imitates the input-output behaviour of the target model, i.e., $S(x)\approx M(x)$. Such surrogates offer a simplified and interpretable approximation of the model’s outputs (Asher et al., 2015; Kudela and Matousek, 2022). However, this type of surrogate does not truly reflect the internal structure (or internal thinking) of the model. To address this limitation, structured surrogate models explicitly accounts for the model’s internal organization (Munk et al., 2022; Páez, 2024). If the target model has a known internal structure:
@@
where $h(x)\in\mathcal{H}$ is the structure, our goal is to construct:
@@
that approximates the internal computational structure of the target model, i.e., $\hat{h}(x)\approx h(x)$, while, at the same time, being faithful to the target model by making highly consistent predictions with it, i.e., $\hat{y}(x)\approx y(x)$. Note that our main goal here is to build a simple surrogate model to imitate a complex model. Improving the ability of a model is not our purpose and is left to future work.
@@ 3.2 Conceptual Framework
A latent argument refers to an internal signal within a model that implicitly conveys supporting or attacking opinions toward a claim. Such signals can arise from different sources, like activations or attention patterns. Because they live in the model’s latent space, these arguments are not directly visible or human-readable, but they still express how intermediate steps “think” about the claim.
@@ 3.3 Symbolic Instantiation for Latent Debate in LLM True-False Prediction
We now describe how the abstract concepts of latent debate can be instantiated in the context of transformer-based LLM true/false prediction tasks (Vaswani et al., 2017). We adopt a symbolic argumentation framework to perform the decision making process, which is transparent and efficient.
Formally, given a query $\mathbf{x}=(x_{1},\ldots,x_{N})$ with a binary label $c\in\{\text{True},\text{False}\}$, a decoder-based LLM generates an answer $\mathbf{y}=(y_{1},\ldots,y_{T})$. Both the query and answer tokens are drawn from the same vocabulary, i.e., $x_{n},y_{t}\in\mathcal{V}$. Each token $y_{t}$ in the answer is generated conditionally based on the preceding tokens and the input query, following the distribution: $y_{t}\sim P(y_{t}\mid\mathbf{y}_{\leq t-1},\mathbf{x})$. In a MAD setting (Du et al., 2024; Liang et al., 2024a), the process involves generating multiple answers $\mathcal{Y}=\{\mathbf{y}^{(1)},\ldots,\mathbf{y}^{(K)}\}$, which may support ($\mathbf{y}^{+}$) or attack ($\mathbf{y}^{-}$) the claim. A final decision is then derived by aggregating these arguments, often through some form of consensus or voting strategies. In contrast, we define that a latent debate takes place inside the model processing claim $\mathbf{x}$ before generating answer $\mathbf{y}$.
3.3.1 Latent Arguments in LLMs
@@
3.3.2 Argument Interpreter in LLMs
To make latent arguments interpretable, the instantiated argument interpreter projects a hidden state into the vocabulary space using the unembedding matrix $\mathbf{W}^{\text{unemb}}\in\mathbb{R}^{|\mathcal{V}|\times d}$. This produces a probability distribution over vocabulary tokens, which has been widely used in mechanistic studies (nostalgebraist, 2020; Belrose et al., 2023). By examining the probabilities assigned to specific tokens True and False, we can quantify the opinion of each latent argument, i.e., how much it supports or attacks the claim.
@@ 3.3.2 Argument Interpreter in LLMs
3.3.3 Thinking Module in LLMs
@@
$\text{cls}(\mathbf{x})$ $=\text{think}(\mathcal{H}),$ (5)
$\mathcal{H}$ $=\left\{\text{interpret}\!\left(\mathbf{h}_{n}^{(l)}\right)\;\middle|\;1\leq n\leq N,\ 1\leq l\leq L-1\right\}$
where $\text{cls}(\mathbf{x})$ is capable of outputting a label associated with the final decision $c$. To perform the thinking step, we adopt a symbolic approach, Quantitative Bipolar Argumentation Framework (QBAF) (Baroni et al., 2019), as the $\textit{think}(\cdot)$ function, which accounts for both supporting and attacking relationships among arguments to yield a coherent, weighted judgment.
@@
Given a set of arguments $\mathcal{A}$, QBAF is capable of considering the overall debate situation and outputting a final strength (as shown in Figure 3(a)), which can be used to obtain the binary (true/false) predictions. In the figure, each node $n_{i}$ corresponds to an argument, and the $\tau(\cdot)$ function assigns the initial strength, indicating its polarity and strength magnitude before propagation. The directed edges between nodes represent the relationships among arguments: edges labeled with - are attacks, indicating that one argument undermines another (two arguments with different polarities), while edges labeled with ++ are supports, meaning that one argument reinforces another. The $\sigma(\cdot)$ beneath the nodes is a function resulting final strengths after applying a chosen gradual semantics, e.g., (Baroni et al., 2015; Rago et al., 2016; Potyka, 2018; Amgoud and Ben-Naim, 2018), which reflects how the collective influence of attackers and supporters modifies the outcome. See Example A.3 in the appendix for a detailed computation process.
Formally, gradual semantics provide the rules that control how initial strengths are updated and result in the computation of final strengths. We define a new semantics. We start from arguments with no attackers or supporters, whose final strength is the same as their initial strength. For the remaining arguments, the final strength is updated along the edges by considering the influence of both attackers and supporters. This process involves two components: aggregation and influence. For an argument $\alpha\!\!\in\!\!\mathcal{A}$, the aggregation step computes its energy $E_{\alpha}\!$ by summing the strengths of its attackers and supporters ($\beta$):